A table -tennis ball is held under water and then released. Described and explain the subsequent motion of the ball


1
Expert's answer
2021-01-04T14:38:25-0500

The mass of the ball is "2.7(g)"


The diameter of the ball is "40 (mm)"


Gravity, buoyant force and the resistance force of the medium act on the ball.


"F_G=mg=0.0027\cdot(-9.8)\approx-0.03(N)"


"F_B=\rho gV=1000\cdot9.8\cdot(4/3)\cdot\pi\cdot0.02^3\approx0.34(N)"


"F_R=-kv".


So, if the force of gravity and the force of resistance of the medium become equal to the modulus of the buoyant force, the ball will move up at a constant speed.